OpenStudy (caerus):

anyone can tell me how to get the roots of this equation ill give medal... x^6+9x^4+24x^2+16=0

OpenStudy (sooobored):

have you tried long dividing or synthetic division factors of 16?

OpenStudy (sooobored):

oh, fyi, make the substitution x^2 =u it will make the wholeee problem much easier

OpenStudy (caerus):

i tried synthetic, and stocked up

OpenStudy (sooobored):

stocked up?

OpenStudy (caerus):

i tried this site and still not get it... http://www.tiger-algebra.com/drill/x~6_9x~4_24x~2_16=0/

OpenStudy (sooobored):

ah ha, cause the answers are all imaginary

OpenStudy (caerus):

that site can do imaginary too

OpenStudy (sooobored):

can it? cause matlab gave me actual imaginary solutions

OpenStudy (caerus):

can you give me the site?

OpenStudy (sooobored):

ok, make the u substituion \[u^3 +9x^2+24u+16\]

OpenStudy (caerus):

cuz this roots, i will work for linear equations w/ constant coefficient

OpenStudy (sooobored):

matlab is a program, a higher tech calculator that engineering students generally use

OpenStudy (sooobored):

ok, note how all positives, so it must have the form (x+n) factors of 16- 1,2,4,8,16 try synthetic division using -1

OpenStudy (caerus):

by that u substitution i get u=-1,-4,-4

OpenStudy (caerus):

what can i do next?

OpenStudy (caerus):

ill substitute the value of u in u=D^2 ?

OpenStudy (sooobored):

no, substitute back in \(x^2\)

OpenStudy (sooobored):

when factored it should look like \( (x^2 +4)(x^2 +4)(x^2 +1)\)

OpenStudy (sooobored):

now, remember how \( x^2-n^2=(x-n)(x+n)\)

OpenStudy (sooobored):

well, \( i *i =-1\)

OpenStudy (caerus):

nice, thank i got the idea of u substitution, thanks for your help

OpenStudy (sooobored):

hence, \(x^2 +n^2= (x-in)(x+in)\)

OpenStudy (sooobored):

yup, no problem